7(x+4)-2{x-3(5+x} =2/3(x-6)
Answer by mananth(16949)


7(x+4)-2{x-3(5+x} =2/3(x-6)
7x+28 -2(x-15-3x) = (2/3)(x-6)
7x +28 -2x+30 +6x= (2x-12)/3
11x+58 =(2x-12)/3
3(11x+58) =2x-12
33x +174=2x-12
31x = -186
x=-186/31
x=-6